# Searching from Area to Point: A Semantic Guided Framework with Geometric Consistency for Accurate Feature Matching

### v5

@@ III-B Geometry Consistency Formulation
In order to leverage the inherent geometry property of area matching to improve matching accuracy and robustness, we proceed to formulate the geometry consistency of area matching. Since the geometry constraint of point matches is formed completely, we utilize the point matches within areas and their fundamental matrices to construct the geometry consistency of area matches. First, the correspondences $\mathcal{P}_{i}=\{(q^{m}_{i},p_{i}^{m})\}_{m}^{M}$ in $\mathcal{A}_{i,\pi(i)}$ can be achieved by $PM$ and the fundamental matrix $F_{i}$ can be calculated as well. Then we form the geometry consistency between $\mathcal{P}_{i}$ and $F_{i}$ by Sampson distance [42]:
@@
$={\sum_{m}^{M}\hat{d}^{m}_{i,i}}=D(F_{i},\mathcal{P}_{i})$
where $(F_{i}q_{i}^{m})_{k}$ represents the k-th entry of the vector $F_{i}q_{i}^{m}$. It should be ideally closed to 0 and reflects the matching precision of $\mathcal{A}_{i,\pi(i)}$, as only the correct area match will produce accurate point matches. Similarly, we can deduce the geometry consistency across area matches. Given two correct area matches $\{\mathcal{A}_{i,\pi(i)},A_{j,\pi(j)}\}$ between the images, they should generate the same Fundamental matrix. Thus the cross Sampson distance ($d_{i,j}$) should be close to 0:
$d_{i,j}$ $=D(F_{i},\mathcal{P}_{j})\rightarrow 0$ (5)
Therefore, in an area match set $\{\mathcal{A}_{i,\pi(i)}\}_{i}^{N}$, assuming most of area matches are correct, the geometry consistency of a specific area match $\mathcal{A}_{i,\pi(i)}$ can be formulated as:
@@
IV Methodology
In this section, we propose the Semantic and Geometry Area Matching method that leverages semantic and geometry priors to construct an efficient feature matching search space, rather than relying on additional point matching [14]. Firstly, we introduce the Semantic Area Matching (Sec. IV-A) to find putative area matches, leveraging semantic prior. Secondly, we describe the Geometry Area Matching (Sec. IV-B), which utilizes geometry consistency to increase the accuracy and robustness of area matching. Finally, the implementation of A2PM framework by combining SGAM with an off-the-shelf detector-free point matcher is illustrated (Sec. IV-C). Symbols in this section are summarised in Tab. I.
IV-A Semantic Area Matching
@@
To establish semantic area matches between images, we propose a detector-based method, including semantic area detection and description, in light of detector-based point matching. Therefore, we first carefully define the semantic area with the goal of enabling it to serve as an efficient search space for feature matching. A typical semantic area is an object-centred area, termed as Semantic Object Area (SOA), whose textured surface or prominent edges favours feature matching. However, some objects (e.g. objects very close to the camera) are so large in the image that the sizes or aspect ratios of the corresponding areas are extremely large, leading to improper search space. Thus, we further propose the Semantic Intersection Area (SIA), which consists of intersecting parts of multiple objects rather than an entire object, to efficiently grab solid features of above large objects for feature matching. Afterwards, we illustrate the matching processes of SOA and SIA (Fig. 3) respectively.
IV-A1 Semantic Object Area Matching
@@
Detecting semantic object areas can be accomplished by identifying the connected components with the object semantic in $I^{s}_{i}$ and establishing their bounding boxes as the boundaries for the areas. For achieving a sparse extraction, we further merge the spatially proximate areas that exhibit the same semantic.
@@
As SOA already contains the object semantic information, its descriptor is intended to distinguish it from other instances of the same semantics in the image. Since close instances are merged by SOA detection, it is crucial to focus on distinguishing the spatially scattered instances, which are likely to have different surroundings. Therefore, we propose the semantic surrounding descriptor to differentiate instances by leveraging semantic information about their surroundings. Initially, four descriptors are constructed as binary zero vectors for each side of SOA. Each bit in these vectors corresponds to a semantic present in the image pair. Subsequently, we record all the semantics that are encountered along the area sides to set the corresponding bits in the binary vectors to true. To enhance scale robustness, we propose the multiscale boundaries capture. By scaling the area boundary to varying degrees, more semantics can be captured by multiscale boundary sides. Finally, the four side vectors are integrated in a specific order to form the SOA descriptor.
@@
The described SOAs are first matched directly by their object semantics, termed as $\{\mathcal{A}_{i,\pi(i)}^{O}\}_{i}$. Subsequently, Hamming distances between the binary descriptors ($\{\mathsf{H}_{i,\pi(i)}\}_{i}$) of initial matches are calculated. Next, matches among areas sharing the same semantic are identified through nearest neighbor search. During this process, if the confidences of two candidate matches for $\alpha_{i}$, termed as $\mathcal{A}_{i,j},\mathcal{A}_{i,k}$, subject to $|\mathsf{H}_{i,j}\!-\!\mathsf{H}_{i,k}|\!<\!T_{da}$, these areas ($\alpha_{i},\beta_{j},\beta_{k}$) are flagged as doubtful. Finally, the matches with $\mathsf{H}_{i,\pi(i)}\!>\!T_{\mathsf{H}}$ are rejected.
@@
3       match inside the area match: $\mathcal{P}_{i}=PM(\mathcal{A}_{i,\pi(i)})$;
@@
Algorithm 1 Geometric Area Match Rejector
IV-A2 Semantic Intersection Area Maching
@@
The detection of SIA involves sliding a window across the $I_{i}^{s}$ to identify areas with abundant semantics. Specifically, the window size is set to the desired area size, and the slide step is half the window size. During the window sliding, areas with more than $3$ different semantics are collected as SIAs. Considering the expensive time cost of window sliding in original $I_{i}^{s}$, we employ a two-layer semantic pyramid. The top layer involves reducing the $I^{s}_{i}$ and window to scale $r$, performing initial detection. The bottom layer is the original $I_{i}^{s}$ , used to further refine the area location. In the refinement, we first calculate the proportion of semantics within each area. To ensure that prominent features remain close to the center of the area, we adjust the area center ($(u,v)$) within the range of area size by minimizing the variance of semantic proportion ($\sigma(u,v)$) within it.
@@
To match SIAs across images, the inside-area semantics is crucial. Therefore, we propose using the semantic proportion, calculated during the detection process, as the SIA descriptor. This descriptor takes the form of a vector, with its length corresponding to the number of semantics present in the image pair. Each element within the vector represents a specific semantic, with its value indicating the proportion of that semantic within the area. To enhance the scale robustness, the descriptors are constructed on multiscale windows and then merged by taking their average. The spatial specificity of the descriptor is improved by partitioning the original area into quarters and constructing descriptors for each sub-area. These sub-area descriptors are then combined to form the final one in a specific order.
@@
Afterwards, SIA matches ($\mathcal{A}^{I}_{i,\pi(i)}$) can be found by nearest neighbour search based on $l_{2}$ distance between descriptors ($\mathsf{L}_{i,\pi(i)}$). Similar to SOA matching, doubtful areas are identified, when $|\mathsf{L}_{i,\pi(i)}\!-\!\mathsf{L}_{j,\pi(j)}|\!<\!T_{da}$ for area matches $\mathcal{A}^{I}_{i,\pi(i)},\mathcal{A}^{I}_{j,\pi(j)}$. Finally, matches with $\mathsf{L}_{i,\pi(i)}\!>\!T_{\mathsf{L}}$ are rejected.
IV-B Geometry Area Matching
@@
Although the SAM is effective in most cases, it tends to overlook local details in images, which can result in semantic ambiguity when multiple instances are present in the image pair. Especially when their semantic surroundings are similar, SAM may obtain doubtful areas and incorrect area matches. Fortunately, area matches, similar to point matches, are inherently constrained by epipolar geometry, which can be utilized to resolve semantic ambiguity. Hence, based on the formulated geometry consistency in Sec. III, we propose Geometry Area Matching (GAM) to refine the results of SAM and fulfill the A2PM framework. GAM incorporates a Predictor (GP, Sec. IV-B1) to identify true matches in doubtful areas, a Rejector (GR, Sec. IV-B2) to eliminate false and inferior area matches and a Global Match Collection module (GMC, Sec. IV-B3) to achieve uniformly distributed point matches.
IV-B1 Geometry Area Match Predictor
@@
The GP aims to determine the true matches among multiple matching possibilities of doubtful areas. Given doubtful areas $\{\alpha_{i}\}_{i}^{H},\{\beta_{i}\}_{i}^{R},R\leq H$ in images $I_{0},I_{1}$ which can not be well matched by SAM, and assume $R$ true area matches exist:
$\mathcal{A}\mathcal{S}_{l}=\{\mathcal{A}_{i,\pi_{l}(i)}\}_{i}^{R}=\{(\alpha_{i},\beta_{\pi_{l}(i)})\}_{i}^{R}$ (7)
@@
where $\mathcal{A}\mathcal{S}_{l}$ is a set of area matches, $\pi_{l}(i)\in[0,R)$ is an index mapping between matched areas with $l$ indicating different area matching possibilities. There are totally $L=\frac{H!}{(H-R)!}$ matching possibilities ($l\in[0,L)$), and only one true area match set ($\mathcal{A}\mathcal{S}_{l^{*}}$) exists with the best geometry consistency, as every area match in $\mathcal{A}\mathcal{S}_{l^{*}}$ is correctly matched. Thus, we can form the geometry consistency of every $\mathcal{A}\mathcal{S}_{l}$ based on Eq. 6:
$G_{\mathcal{A}\mathcal{S}_{l}}=\frac{1}{R}\sum_{i}^{R}G_{\mathcal{A}_{i,\pi_{l}(i)}}$ (8)
which can be further written in probabilistic form:
$P(\mathcal{A}\mathcal{S}_{l})=\exp(-G_{\mathcal{A}\mathcal{S}_{l}})$ (9)
Therefore, the $\mathcal{A}\mathcal{S}_{l^{*}}$ can be achieved by geometry consistency maximization:
[…]
@@
IV-B2 Geometry Area Match Rejector
Following the prediction, the GR leverages geometry consistency to identify and disregard potential false matches, thereby enhancing the accuracy of the area matching. Given an area match set $As=\{\mathcal{A}_{i,\pi(i)}\}_{i}^{S}$ achieved by SAM and GP, the geometry consistency of each $\mathcal{A}_{i,\pi(i)}$ can be measured by $G_{\mathcal{A}_{i,\pi(i)}}$ (Eq. 6). Then, matches with $G_{\mathcal{A}_{i,\pi(i)}}$ exceeding a specific threshold can be discarded as false or inaccurate. In practice, the threshold $T_{GR}$ is based on the mean self-geometry consistency (Eq. 4) with a weight $\phi$. The Transformer-based matcher [10, 37] is embedded in GR to acquire precise point matches. The specific process is illustrated in Algorithm 1.
@@ IV-B2 Geometry Area Match Rejector
3if $SP_{\{\mathcal{A}_{i,\pi(i)}\}_{i}^{T}}<T_{SP}$  then
4       initialize the empty collected match list: ${\mathcal{P}}_{g}=\varnothing$;
5       calculate the fundamental matrix $F_{a}$ and mean Sampson distance $\frac{1}{T}\sum_{i}^{T}d_{a,i}$ of inside-area point matches $\{\mathcal{P}_{i}\}_{i}^{T}$;
@@
8             get the single match Sampson distance $\hat{d}^{m}_{a,g}$;
9             if $\hat{d}^{m}_{a,g}<=\frac{1}{T}\sum_{i}^{T}d_{a,i}$ then collect the match $(p^{m}_{g},q^{m}_{g})$ into ${\mathcal{P}}_{g}$;
@@ 12 Output the collected global matches: ${\mathcal{P}}_{g}$;
IV-B3 Global Match Collection
The precision of point matches within area matches is remarkable, termed as $\{\mathcal{P}_{i},\mathcal{A}_{i,\pi(i)}\}_{i}^{T}$, as a result of the well-defined search spaces achieved through both semantic prior and geometric consistency. However, the distribution of these matches depends on the specific scenes. If less semantic information is available in the scene, there will be a small number of area matches. Consequently, the point matches will cluster, which has a negative impact on the downstream tasks [42]. To enhance the robustness against scenes with limited semantic information, we propose the GMC module to collect global matches ($\mathcal{P}_{g}$) utilizing the geometry constraint of accurate inside area matches, when few area matches exist (determined by size proportion threshold $T_{SP}$, which represents the proportion threshold of the image occupied by matched areas in the image pair). The detailed algorithm is presented in Algorithm 2.
IV-C Framework Implementation
In summary, the A2PM framework follows the steps outlined below. First, the framework obtains the putative area matches ($\{A^{*}_{i,\pi(i)}\}_{i}^{K}$) and the doubtful areas ($\{\alpha_{i}\}_{i}^{H}$, $\{\beta_{i}\}_{i}^{R}$, where $R\leq H$) between the images through semantic segmentation ($I^{s}_{0},I^{s}_{1}$) using the SAM algorithm:
$\small\{A^{*}_{i,\pi(i)}\}_{i}^{K},\{\alpha_{i}\}_{i}^{H},\{\beta_{i}\}_{i}^{R}=SAM(I^{s}_{0},I^{s}_{1})$ (11)
Next, the doubtful areas are cropped from original images ($I_{0},I_{1}$) and matched by an off-the-shelf detector-free Point Matcher ($PM$) integrated in GP to achieve inside-area correspondences for geometry consistency calculation. Then, the $R$ true area matches can be identified by $GP_{PM}$:
$\{A^{*}_{i,\pi(i)}\}_{i}^{R}=GP_{PM}(\{\alpha_{i}\}_{i}^{H},\{\beta_{i}\}_{i}^{R},I_{0},I_{1})$ (12)
Afterwards, the accurate area and point matches inside them are achieved by $GR$ integrated with $PM$ ($T\leq K+R$):
$\{\mathcal{A}_{i,\pi(i)},\mathcal{P}_{i}\}_{i}^{T}=GR_{PM}(\{A^{*}_{i,\pi(i)}\}_{i}^{K+R},I_{0},I_{1})$ (13)

### v6

@@ III-A2 Geometry Consistency of Area Matching
In order to leverage the intrinsic geometry property to improve matching accuracy and robustness, we proceed to formulate the geometry consistency of area matching. Since point-level geometry constraint is formed completely [40], we utilize the epipolar geometry of point matches within areas to construct the area-level geometry consistency. First, the correspondences $\mathcal{P}_{i}=\{(q^{m}_{i},p_{i}^{m})\}_{m}^{M}$ in $\mathcal{A}_{i,\pi(i)}$ can be achieved by $PM$ and the fundamental matrix $F_{i}$ can be calculated as well. Then we form the geometry consistency between $\mathcal{P}_{i}$ and $F_{i}$ by Sampson distance [40]:
@@
$={\sum_{m}^{M}\hat{d}^{m}_{i,i}}=D(F_{i},\mathcal{P}_{i}),$
where $(F_{i}q_{i}^{m})_{k}$ represents the k-th entry of the vector $F_{i}q_{i}^{m}$. It should be ideally close to 0 and reflects the matching precision of $\mathcal{A}_{i,\pi(i)}$, as only the correct area match produces accurate point matches. Similarly, we can infer the geometry consistency across area matches. For two correct area matches $\{\mathcal{A}_{i,\pi(i)},A_{j,\pi(j)}\}$ between the images, they should ideally yield the same fundamental matrix. Thus the cross Sampson distance ($d_{i,j}$) should be close to 0:
$d_{i,j}$ $=D(F_{i},\mathcal{P}_{j})\rightarrow 0.$ (5)
Therefore, within an area match set $\{\mathcal{A}_{i,\pi(i)}\}_{i}^{N}$, assuming the most of area matches are correct, the geometry consistency of a specific area match $\mathcal{A}_{i,\pi(i)}$ can be formulated as:
@@
III-B Semantic Area Matching
Figure 3: Semantic Area Matching (SAM). Two types of semantic areas are proposed by SAM. Both of them are detected and described by hand-crafted semantic feature. Then area matches are established by nearest neighbour search based on descriptor distance.
@@ III-B Semantic Area Matching
To find semantic area matches between images, we propose Semantic Area Matching (SAM), which adopts a detection-and-description manner similar to sparse point matching [7]. Particularly, we first propose two typical semantic areas with the goal of achieving a better integration between semantics and search space. The first area is an object-centric area, termed as Semantic Object Area (SOA), where the textured surface and prominent edges of the inside object favour point matching. However, some objects (e.g. objects very close to the camera) are so large in the image that the sizes or aspect ratios of the corresponding areas are extremely large, leading to improper search space. Thus, we further propose the Semantic Intersection Area (SIA), which consists of intersecting parts of multiple objects rather than an entire object, to efficiently grab solid features in the intersection of above large objects. Afterwards, we illustrate the detection-and-description matching processes of SOA and SIA (Fig. 3). Given semantic segmentation of images ($I^{s}_{i}$), both of them leverage hand-crafted semantic features to achieve area matches with sufficient accuracy, benefiting from semantic robustness in matching.
III-B1 Matching of SOA
@@
Detecting semantic object areas can be accomplished by identifying the connected components with the object semantic in $I^{s}_{i}$ and utilizing their bounding boxes as the boundaries for the areas. For achieving a sparse extraction, we merge the spatially proximate areas that exhibit the same semantic.
@@
As SOA already contains the distinguishable object semantic information, its descriptor is designed to differentiate it from other instances of the same semantics in the image. Since close instances are merged in detection, it is crucial to focus on distinguishing the spatially scattered instances, which are likely to have different surroundings. Therefore, we propose the semantic surrounding descriptor to differentiate instances by leveraging semantic information about their surroundings. Similar to BRIEF [41] descriptor in point matching, this descriptor is also a binary vector where each bit corresponds to a semantic present in the image pair. A bit with a value of $1$ indicates the presence of the semantic along the area sides, while a value of 0 indicates its absence, thereby representing the semantic context surrounding the area. To enhance scale robustness, we propose the multiscale boundaries capture. By scaling the area boundary to varying degrees, more semantics can be captured (see Fig. 3 top ‘Area Description’ part.).
@@
The SOAs are first matched directly by their inside object semantics. Subsequently, Hamming distances [41] between the descriptors of initial matches are calculated. Next, area matches sharing the same semantic are determined through the nearest neighbor search. During this process, matches with descriptor distances larger than the threshold $T_{H}$ are rejected. Some SOAs may be labeled as doubtful, due to similar descriptor distances (identified by threshold $T_{da}$) of multiple match candidates. These doubtful areas will be handled in the geometric area matching.
@@
3       perform point matching inside the area match: $\mathcal{P}_{i}=PM(\mathcal{A}_{i,\pi(i)})$;
@@
Algorithm 1 Geometric Area Matching Rejector
III-B2 Matching of SIA
@@
The detection of SIA involves sliding a window across the $I_{i}^{s}$ to identify areas with abundant semantic specificity. In particular, the window size is set to the desired area size, and the slide step is half the window size. During the window sliding, areas with more than $3$ different semantics are collected as SIAs. Considering the expensive time cost of window sliding in original $I_{i}^{s}$, we employ a two-layer semantic pyramid, similar to classical image pyramid. The top layer involves reducing the $I^{s}_{i}$ and window to scale $r$, performing initial detection. The bottom layer is the original $I_{i}^{s}$ , used to further refine the area location. In the refinement, we first calculate the proportion of every semantics within each area. Then, we adjust the center of each area within a certain range (the area size) to ensure uniform proportion of different semantics in the area, by minimizing the variance of semantic proportion ($\sigma(u,v)$) within it (see Fig. 3 bottom ‘Area Detection’ part.).
@@
The inside-area semantics are crucial to match SIAs. Therefore, we propose utilizing the semantic proportion, calculated during the detection process, as the SIA descriptor. Similar to SOA, this descriptor is also a vector, but it is not binary. Each element of the vector represents a specific semantic, with its value indicating the proportion of that semantic within the area. To enhance the scale robustness, the descriptors are constructed on multiscale windows with the same center and then merged by taking their average.
@@
Afterwards, SIA matches can be found by nearest neighbour search based on $l_{2}$ distance between descriptors. Similar to SOA matching, doubtful areas are identified by $T_{da}$ and inferior matches are rejected by $T_{l}$.
III-C Geometry Area Matching
@@
Although the SAM is effective in most cases, it tends to overlook local details in images, potentially leading to semantic ambiguity when multiple instances are present in the image pair. Especially when semantic surroundings of instances are also similar, SAM may obtain doubtful areas and incorrect area matches. Fortunately, area matches, similar to point matches, are inherently constrained by epipolar geometry, which can be utilized to resolve the semantic ambiguity. Hence, based on the formulated geometry consistency in Sec. III-A, we propose Geometry Area Matching (GAM) to refine the results of SAM and fulfill the A2PM framework. GAM incorporates a Predictor (GP, Sec. III-C1) to identify true matches in doubtful areas, a Rejector (GR, Sec. III-C2) to eliminate inferior area matches and a Global Match Collection module (GMC, Sec. III-C3) to achieve uniformly distributed matches.
III-C1 Geometry Area Matching Predictor
@@
The GP aims to determine the true matches among multiple matching possibilities of doubtful areas. Given doubtful areas $\{\alpha_{i}\}_{i}^{H},\{\beta_{j}\}_{j}^{R},R\leq H$ in $I_{0},I_{1}$ which can not be confidently matched by SAM, and assume $R$ correct area matches exist:
$\mathcal{A}\mathcal{S}_{l}=\{\mathcal{A}_{i,\pi_{l}(i)}\}_{i}^{R}=\{(\alpha_{i},\beta_{\pi_{l}(i)})\}_{i}^{R},$ (7)
@@
where $\mathcal{A}\mathcal{S}_{l}$ is a set of area matches, $\pi_{l}(i)\in[0,R)$ is an index mapping between matched areas with $l$ indicating different area matching possibilities. There are totally $L={H!}/{(H-R)!}$ matching possibilities ($l\in[0,L)$), and only one true area match set ($\mathcal{A}\mathcal{S}_{l^{*}}$) exists with the best geometry consistency, where every area match is correctly matched. Thus, we first form the geometry consistency of any $\mathcal{A}\mathcal{S}_{l}$ based on Eq. 6:
$G_{\mathcal{A}\mathcal{S}_{l}}=\frac{1}{R}\sum_{i}^{R}G_{\mathcal{A}_{i,\pi_{l}(i)}}.$ (8)
Then the likelihood of $\mathcal{A}\mathcal{S}_{l}$ can be represented as:
$P(\mathcal{A}\mathcal{S}_{l})=\exp(-G_{\mathcal{A}\mathcal{S}_{l}}).$ (9)
Therefore, the true match set $\mathcal{A}\mathcal{S}_{l^{*}}$ can be achieved by likelihood maximization:
${As}_{{l}^{*}}=\arg\max_{l}~{}P(\mathcal{A}\mathcal{S}_{l}).$ (10)
@@
III-C2 Geometry Area Matching Rejector
Following the prediction, the GR leverages geometry consistency to identify and reject potential false matches, thereby enhancing the accuracy of the area matching. Given an area match set $\mathcal{A}\mathcal{S}=\{\mathcal{A}_{i,\pi(i)}\}_{i}^{S}$ achieved by SAM and GP, the geometry consistency of each $\mathcal{A}_{i,\pi(i)}$ can be measured by $G_{\mathcal{A}_{i,\pi(i)}}$ (Eq. 6). Then, matches with $G_{\mathcal{A}_{i,\pi(i)}}$ exceeding a specific threshold $T_{GR}$ can be discarded as inaccurate. In practice, the $T_{GR}$ is based on the mean self-geometry consistency (Eq. 4) with a weight $\phi$. The point matcher is embedded in GR to acquire point matches. The specific process is illustrated in Algorithm 1.
@@ III-C2 Geometry Area Matching Rejector
3initialize the ${\mathcal{P}}_{g}=\varnothing$
4if $SP_{\{\mathcal{A}_{i,\pi(i)}\}_{i}^{T}}<T_{SP}$  then
5       calculate the fundamental matrix $F_{a}$ and mean Sampson distance $\frac{1}{T}\sum_{i}^{T}d_{a,i}$ (by Eq. 5) of inside-area point matches $\{\mathcal{P}_{i}\}_{i}^{T}$;
@@
8             get the single match Sampson distance ${d}^{m}_{a,g}$;
9             if ${d}^{m}_{a,g}<=\frac{1}{T}\sum_{i}^{T}d_{a,i}$ then collect the match $(p^{m}_{g},q^{m}_{g})$ into ${\mathcal{P}}_{g}$;
@@ 12 Output the collected global matches: ${\mathcal{P}}_{g}$;
III-C3 Global Match Collection
The precision of point matches within area matches is guaranteed, termed as $\{\mathcal{P}_{i},\mathcal{A}_{i,\pi(i)}\}_{i}^{T}$, as a result of the improved search spaces achieved through both semantic prior and geometric consistency. However, the distribution of these matches depends on the specific scenes. If less semantic information is available in the scene, there will be a small number of area matches. Consequently, the point matches will cluster, which has a negative impact on the downstream tasks [40]. To enhance the robustness against scenes with limited semantic information, we propose the Global Match Collection (GMC) module to collect global matches ($\mathcal{P}_{g}$) utilizing the geometry constraint of accurate inside-area point matches in scenes with less semantic content. These scenes are identified based on a size proportion threshold $T_{SP}$, which is the proportion threshold of the image occupied by matched areas in the image pair. The detailed algorithm is presented in Algorithm 2.
III-D Framework Implementation
The overall A2PM framework follows the steps outlined below. Firstly, given the semantic segmentation of the input image pair ($I^{s}_{0},I^{s}_{1}$), the framework obtains putative area matches ($\{\mathcal{A}^{*}_{i,\pi(i)}\}_{i}^{K}$) and doubtful areas ($\{\alpha_{i}\}_{i}^{H}$, $\{\beta_{i}\}_{i}^{R}$, $R\leq H$) between the images using the SAM algorithm:
$\small\{\mathcal{A}^{*}_{i,\pi(i)}\}_{i}^{K},\{\alpha_{i}\}_{i}^{H},\{\beta_{i}\}_{i}^{R}=SAM(I^{s}_{0},I^{s}_{1}).$ (11)
Next, the doubtful areas are cropped from original images ($I_{0},I_{1}$) and matched by a Point Matcher ($PM$) integrated in $GP_{PM}$ to achieve inside-area correspondences for geometry consistency calculation. Then, the true area matches ($\{\mathcal{A}^{*}_{i,\pi(i)}\}_{i}^{R}$) can be identified by $GP_{PM}$:
$\{\mathcal{A}^{*}_{i,\pi(i)}\}_{i}^{R}=GP_{PM}(\{\alpha_{i}\}_{i}^{H},\{\beta_{i}\}_{i}^{R},I_{0},I_{1}).$ (12)
Afterwards, the accurate area and point matches inside them ($\{\mathcal{A}_{i,\pi(i)},\mathcal{P}_{i}\}_{i}^{T}$) are achieved by $GR_{PM}$ integrated with $PM$ (The areas are cropped and subsequently matched by PM.):
$\{\mathcal{A}_{i,\pi(i)},\mathcal{P}_{i}\}_{i}^{T}=GR_{PM}(\{A^{*}_{i,\pi(i)}\}_{i}^{K+R},I_{0},I_{1}).$ (13)